Macawi, who weighs 108 pounds, is seated 5 feet from the fulcrum of a seesaw. Barbara is seated on the same side of the seesaw, 2 feet farther from the fulcrum than Macawi. Barbara weighs 96 pounds. The seesaw is balanced when Sue, who weighs 101 pounds, sits on the other side. How far is Sue from the fulcrum?
Help would be greatly appreciated thank you :)

Answer by ikleyn(52878)

.
Macawi, who weighs 108 pounds, is seated 5 feet from the fulcrum of a seesaw. Barbara is seated on the same side of the seesaw,
2 feet farther from the fulcrum than Macawi. Barbara weighs 96 pounds. The seesaw is balanced when Sue,
who weighs 101 pounds, sits on the other side. How far is Sue from the fulcrum?
Help would be greatly appreciated thank you :)
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

The condition of equilibrium is equality of force moments (see  this site,  for example,

as well as the lesson Using proportions to solve word problems in Physics in this site),

This condition takes the form of the equation

108*5 + 96*7 = 101*x,

where 7 = 5 + 2 and x is the Sue position.

Solve it for x, and you will get the answer.